On October 1,2021, Venus Corp owns 15,000 fair value through other comprehensive income share acquired at a cost of P 345,000. The shares represent 15% of the shares outstanding of Mercury Corporation. On the same date, Mercury Corp, declared P 8 cash dividends on its outstanding shares payable to stock holder on October 31. However, on October 31, Mercury Corp issued 1 share for every 5 shares held by the shareholders in lieu of the supposed cash dividends previously declared.

Prepare all the necessary entries on


a. October 1, 2021

b. October 31, 2021



1
Expert's answer
2021-06-03T18:39:57-0400

Entries on:

(a) October 1, 2021


In the books of Venus corp:

Dividends accounts (Dr.) P (15,000 "\times" 8) = P 120,000

Subsidiary (Mercury Corp/Investment) account (Cr.) by P 120,000


(b) October 31, 2021:


After right issue:

Subsidiary account (Dr.) by P ["\frac{15,000}{5}" ] "\times" 8 = P 120,000

Share Capital account (Cr.) by P 120,000